OpenStudy (anonymous):

Prove directly: Given that a, b, c is a Pythagorean triple, prove that qa, qb, and qc is a Pythagorean triple where q is contained in R, q does not equal 0.

OpenStudy (anonymous):

\[ (qb)^2 + (qc)^2 = q^2 b^2 + q^2 c^2= q^2(b^2+c^2) \] Can you finish it now?

OpenStudy (anonymous):

Notice that \[ b^2 + c^2 = a^2 \]

OpenStudy (anonymous):

\[ (qb)^2 + (qc)^2 = q^2 b^2 + q^2 c^2= q^2(b^2+c^2)= q^2 a^2 \]

OpenStudy (anonymous):

\[ q^2 a^2 = (qa)^2 \]

OpenStudy (anonymous):

Can you explain in more detail from the first step.

OpenStudy (anonymous):

Do you know that \[ (x y)^2= x^2 y^2 \]?

OpenStudy (anonymous):

Do you know that \( mn + m p = m(n+p) \)?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

In this case, you should understand step 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you understand the whole process now?

OpenStudy (anonymous):

Can you write it step by step. Im losing the beginning.